Solve for Y(s), the Laplace transform of the solution y(t) to the initial value problem below. y + 8y = 2 - 6, y(0) = 0, y
Answer #1

y+8y = t2-6. Let Lly (W) = Y(S) y (0)=0, Y!(0)= -2 Noul taking Laplace tansform both sides, oleman bla le ala -- Y (8) tt 3